OpenStudy (anonymous):

The work of a student to solve a set of equations is shown: Equation A: y = 4 – 2z Equation B: 4y = 2 – 4z Step 1: –4(y) = –4(4 – 2z) [Equation A is multiplied by –4.] 4y = 2 – 4z [Equation B] Step 2: –4y = 4 – 2z [Equation A in Step 1 is simplified.] 4y = 2 – 4z [Equation B] Step 3: 0 = 6 – 6z [Equations in Step 2 are added.] Step 4: 6z = 6 Step 5: z = 1 In which step did the student first make an error? (5 points) Step 1 Step 3 Step 4 Step 2

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

Give me a sec please

OpenStudy (godlovesme):

And can u screenshot it :)

OpenStudy (anonymous):

idk how

OpenStudy (godlovesme):

http://www.take-a-screenshot.org

OpenStudy (anonymous):

The work of a student to solve a set of equations is shown: Equation A: y = 4 – 2z Equation B: 4y = 2 – 4z Step 1: –4(y) = –4(4 – 2z) [Equation A is multiplied by –4.] 4y = 2 – 4z [Equation B] Step 2: –4y = 4 – 2z [Equation A in Step 1 is simplified.] 4y = 2 – 4z [Equation B] Step 3: 0 = 6 – 6z [Equations in Step 2 are added.] Step 4: 6z = 6 Step 5: z = 1

OpenStudy (godlovesme):

ok nvm what do u think the answer is?

OpenStudy (anonymous):

step 2

OpenStudy (godlovesme):

sorry i was trying to find the right answer but i couldnt i guess it's 2 not sure thou :/ sorry

OpenStudy (anonymous):

thank you

OpenStudy (godlovesme):

no problem :)